OpenStudy (anonymous):

Translate this word problem as a system of equations and then solve using substitution. The sum of two numbers is 84. One number is three times the other. Find the numbers.

Parth (parthkohli):

Well let's take them as x and y. \(\Large \color{Black}{\Rightarrow } x + y = 84\) \(\Large \color{Black}{\Rightarrow x = 3y}\)

OpenStudy (anonymous):

let first number is x second number is y the sum of two numbers is 84 so, x+y=84---- a x is 3y x=3y now put x=3y in eq a 3y+y=84 4y=84 y=84/4 y=21 now as x=3y x=3* 21 x= 63